OpenStudy (anonymous):

Alternating Series. Determine if Sigma (n=1 to infinity) of (-1)^n (2n/15n+1) converges. I know that we take the limit of (2n/15n+1) as n approaches infinity. But; I'm not sure what to do since we get infinity/infinity when we take the limit...

OpenStudy (amistre64):

convergence looks at the ration of next/now to see how fast it goes to zero

OpenStudy (amistre64):

but it helps if a read it right

OpenStudy (amistre64):

we can determine this from the horizontal asymptote i believe

OpenStudy (amistre64):

2n/(15n+1) , for larger and larger value of n looks more like: 2n/15n would you agree?

OpenStudy (anonymous):

Absolutely! !!

OpenStudy (anonymous):

So now we have an (An) and (bn) to try the limit comparison test?

OpenStudy (amistre64):

then for large value of n. this limits out to 2/15

OpenStudy (amistre64):

we dont even need to do a limit comparison test since this thing doesnt even go to zero

OpenStudy (amistre64):

if the limit of the function doesnt go to zero, it diverges

OpenStudy (anonymous):

Oooo OK!

OpenStudy (anonymous):

Could we also use L'Hopital's rule since we get an indeterminate form of infinity/infinity? Which would then give us that 2/15?

OpenStudy (amistre64):

no need, but yes, you could if you got bored enough to carry it on :)

OpenStudy (anonymous):

YES!

OpenStudy (anonymous):

I think it might be a little different since it is an alternating sequence. Consider this.\[\sum_{n=1}^{\infty}(-1)^n (2n/(15n+1))=\sum_{k=1}^{\infty}(-\frac{4k-2}{30k-14}+\frac{4k}{30k+1})\]

OpenStudy (anonymous):

The first term is the 2k-1 term, corresponding with the odd values of n (which are all negative), while the second term is the 2k term that corresponds to the even values of n.

OpenStudy (anonymous):

Continuing....\[=\sum_{k=1}^{\infty})\frac{(2-4k)(30k+1)+4k(30k-14)}{(30k-14)(30k+1)}\]

OpenStudy (anonymous):

\[=\sum_{k=1}^{\infty}\frac{-120k^2+56k+2+120k-56k}{900k^2-390k-14}=\sum_{k=1}^{\infty}\frac{2}{900k^2-390k-14}\]

OpenStudy (anonymous):

Pretty sure that converges.

OpenStudy (anonymous):

Certainly, if the series is not alternating, it diverges, but since it alternates it will converge.

OpenStudy (anonymous):

@squirefreak Take a look at my work on your problem. I came to a different conclusion.